# Can Direct Latent Model Learning Solve Linear Quadratic Gaussian Control?

### v2

@@ 4 Theoretical guarantees
Theorem 1 below offers a finite-sample guarantee for our approach. Overall, it confirms cost-driven latent model learning (Algorithm 1) as a viable path to solving LQG control.
@@
Given an unknown LQG system (2.1), under Assumptions 1234 and 5, if we run Algorithm 1 with $n\geq\mathrm{poly}(T,d_{x},d_{y},d_{u},\log(1/p))$, then with probability at least $1-p$, state representation function $(\hat{M}_{t})_{t=0}^{T}$ is $\mathrm{poly}(\ell,d_{x},d_{y},d_{u})n^{-1/4}$ optimal in the first $\ell$ steps, and $\mathrm{poly}(\nu^{-1},T,d_{x},d_{y},d_{u})n^{-1/2}$ optimal in the next $(T-\ell)$ steps. Also, the learned controller $(\hat{K}_{t})_{t=0}^{T-1}$ is $\mathrm{poly}(\ell,m,d_{x},d_{y},d_{u})c^{\ell}n^{-1/4}$ optimal for some dimension-free constant $c>0$ depending on system parameters in the first $\ell$ steps, and $\mathrm{poly}(T,m,d_{x},d_{y},d_{u},\log(1/p))n^{-1}$ optimal in the last $(T-\ell)$ steps.
@@
From Theorem 1, we observe a separation of the sample complexities before and after time step $\ell$, resulting from the loss of the full-rankness of $(\mathrm{Cov}(z^{\ast}_{t}))_{t=0}^{\ell-1}$ and $(M^{\ast}_{t})_{t=0}^{\ell-1}$. In more detail, a proof sketch goes as follows. Quadratic regression guarantees that $\hat{N}_{t}$ converges to $N^{\ast}_{t}$ at a rate of $n^{-1/2}$ for all $0\leq t\leq T$. Before step $\ell$, $\hat{M}_{t}$ suffers a square root decay of the rate to $n^{-1/4}$ because $M^{\ast}_{t}$ may not have rank $d_{x}$. Since $(\hat{z}_{t})_{t=0}^{\ell-1}$ may not have full-rank covariances, $(A^{\ast}_{t})_{t=0}^{\ell-1}$ are only recovered partially. As a result, $(\hat{K}_{t})_{t=0}^{\ell-1}$ may not stabilize $(A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{\ell-1}$, causing the exponential dependence on $\ell$. This means if $n$ is not big enough, this controller may be inferior to zero control, since the system $(A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{\ell-1}$ is uniformly exponential stable (Assumption 1) and zero control has suboptimality gap linear in $\ell$. After step $\ell$, $\hat{M}_{t}$ retains the $n^{-1/2}$ sample complexity, and so do $(\hat{A}_{t},\hat{B}_{t})$; the certainty equivalent controller then has an order of $n^{-1}$ suboptimality gap for linear quadratic control (Mania et al., 2019).
Theorem 1 states the guarantees for the state representation function $(\hat{M}_{t})_{t=0}^{T}$ and the controller $(\hat{K}_{t})_{t=0}^{T-1}$ separately. One may wonder the suboptimality gap of $\hat{\pi}=(\hat{M}_{t},\hat{K}_{t})_{t=0}^{T-1}$ in combination; after all, this is the output policy. The new challenge is that a suboptimal controller is applied to a suboptimal state estimation. An exact analysis requires more effort, but a reasonable conjecture is that $(\hat{M}_{t},\hat{K}_{t})_{t=0}^{T-1}$ has the same order of suboptimality gap as $(\hat{K}_{t})_{t=0}^{T-1}$: before step $\ell$, the extra suboptimality gap resulted from $(\hat{M}_{t})_{t=0}^{\ell-1}$ can be analyzed by considering perturbation $\hat{K}_{t}(\hat{M}_{t}-M^{\ast}_{t})h_{t}$ on controls; after step $\ell$, similar to the analysis of the LQG suboptimality gap in (Mania et al., 2019), the overall suboptimality gap can be analyzed by a Taylor expansion of the value function at $(M^{\ast}_{t},K^{\ast}_{t})_{t=\ell}^{T-1}$, with $(\hat{K}_{t}\hat{M}_{t}-K^{\ast}_{t}M^{\ast}_{t})_{t=\ell}^{T-1}$ being perturbations.
4.1 Proof of Theorem 1
@@
Recall that Algorithm 1 has three main steps: state representation learning (CoReL, Algorithm 2), latent system identification (SysId, Algorithm 3), and planning by RDE (2.4). Correspondingly, the analysis below is organized around these three steps.
Recovery of the state representation function. By Proposition 3, with $u_{t}=\mathcal{N}(0,\sigma_{u}^{2}I)$ for all $10\leq t\leq T-$ 1 to system (2.1), the $k$-step cumulative cost starting from step $t$, where $k=1$ for $10\leq t\leq\ell-$ 1 and $k=m\wedge(T-t+1)$ for $\ell\leq t\leq T$, is given under the normalized parameterization by
Let $N^{\ast}_{t}:=(M^{\ast}_{t})^{\top}M^{\ast}_{t}$ for all $0\leq t\leq T$. For quadratic regression, Lemma 2 (detailed later) and the union bound over all time steps guarantees that as long as $n\geq aT^{4}(d_{y}+d_{u})^{4}\log(aT^{3}(d_{y}+d_{u})^{2}/p)\log(T/p)$ for an absolute constant $a>0$, with probability at least $1-p$, for all $0\leq t\leq T$,
$\|\hat{N}_{t}-N^{\ast}_{t}\|_{F}=\mathcal{O}(kt^{2}(d_{y}+d_{u})^{2}d_{x}^{1/2}n^{-1/2}\log^{1/2}(1/p)).$
$\|\widetilde{N}_{t}-N^{\ast}_{t}\|_{F}\leq\|\hat{N}_{t}-N^{\ast}_{t}\|_{F}.$
The low-rank factorization is essentially a combination of low-rank approximation and matrix factorization. For $d_{h}<d_{x}$, $\widetilde{M}_{t}=:[\Sigma^{1/2}U^{\top};0_{(d_{x}-d_{h})\times d_{h}}]$ constructed by padding zeros satisfies $\widetilde{M}_{t}^{\top}\widetilde{M}_{t}=\widetilde{N}_{t}$. For $d_{h}\geq d_{x}$, construct $\widetilde{M}_{t}=:\Sigma_{d_{x}}^{1/2}U_{d_{x}}^{\top}$, where $\Sigma_{d_{x}}$ is the left-top $d_{x}\times d_{x}$ block in $\Sigma$ and $U_{d_{x}}$ consists of $d_{x}$ columns of $U$ from the left. By the Eckart-Young-Mirsky theorem, $\widetilde{M}_{t}^{\top}\widetilde{M}_{t}=U_{d_{x}}\Sigma_{d_{x}}U_{d_{x}}^{\top}$ satisfies
Hence,
$\|\widetilde{M}_{t}^{\top}\widetilde{M}_{t}-N^{\ast}_{t}\|_{F}\leq\;$ $\|\widetilde{M}_{t}^{\top}\widetilde{M}_{t}-\widetilde{N}_{t}\|_{F}+\|\widetilde{N}_{t}-N^{\ast}_{t}\|_{F}$
[…]
=\;= $\mathcal{O}(kt^{2}(d_{y}+d_{u})^{2}d_{x}^{1/2}n^{-1/2}\log^{1/2}(1/p)).$
From now on, we consider $10\leq t\leq\ell-$ 1 and $\ell\leq t\leq T$ separately, since, as we will show, in the latter case we have the additional condition that $\mathrm{rank}(M^{\ast}_{t})=d_{x}$.
For $10\leq t\leq\ell-$ 1, $k=1$. By Lemma 4 to be proved later, there exists a $d_{x}\times d_{x}$ orthogonal matrix $S_{t}$, such that $\|\widetilde{M}_{t}-S_{t}M^{\ast}_{t}\|_{2}\leq\|\widetilde{M}_{t}-S_{t}M^{\ast}_{t}\|_{F}=\mathcal{O}(t(d_{y}+d_{u})d_{x}^{3/4}n^{-1/4}\log^{1/4}(1/p))$. Recall that $\hat{M}_{t}=\textsc{TruncSV}{}(\widetilde{M}_{t},\theta)$; that is, $\hat{M}_{t}=(\mathbb{I}_{[\theta,+\infty)}(\Sigma_{d_{x}}^{1/2})\odot\Sigma_{d_{x}}^{1/2})U_{d_{x}}^{\top}$. Then
$\|\hat{M}_{t}-\widetilde{M}_{t}\|_{2}=\;$ $\|(\mathbb{I}_{[\theta,+\infty)}(\Sigma_{d_{x}}^{1/2})\odot\Sigma_{d_{x}}^{1/2}-\Sigma_{d_{x}}^{1/2})U_{d_{x}}^{\top}\|_{2}$
[…]
Hence, the distance between $\hat{M}_{t}$ and $M^{\ast}_{t}$ satisfies
[…]
$\theta$ should be chosen in such a way that it keeps the error on the same order; that is, $\theta=\mathcal{O}(t(d_{y}+d_{u})d_{x}^{3/4}n^{-1/4}\log^{1/4}(1/p))$. As a result, since $\hat{z}_{t}=\hat{M}_{t}h_{t}$ and $z^{\ast}_{t}=M^{\ast}_{t}h_{t}$,
$\|\hat{z}_{t}-S_{t}z^{\ast}_{t}\|=\;$ $\|(\hat{M}_{t}-S_{t}M^{\ast}_{t})h_{t}\|\leq\|\hat{M}_{t}-S_{t}M^{\ast}_{t}\|_{2}\|h_{t}\|.$
Since $h_{t}=[y_{0:t};u_{0:(t-1)}]$ whose $\ell_{2}$-norm is sub-Gaussian with its mean and sub-Gaussian norm bounded by $\mathcal{O}((t(d_{y}+d_{u}))^{1/2})$, $\|\hat{z}_{t}-S_{t}z^{\ast}_{t}\|$ is sub-Gaussian with its mean and sub-Gaussian norm bounded by
$\mathcal{O}((t(d_{y}+d_{u}))^{3/2}d_{x}^{3/4}n^{-1/4}\log^{1/4}(1/p)).$
On the other hand, threshold $\theta$ ensures a lower bound on $\sigma_{\min}^{+}(\sum\nolimits_{i=1}^{n}z_{t}^{(i)}(z_{t}^{(i)})^{\top})$. As shown in the proof of Lemma 5, this property is important for ensuring the system identification outputs $\hat{A}_{t}$ and $\hat{B}_{t}$ have bounded norms. Specifically,
$\sigma_{\min}^{+}(\sum\nolimits_{i=1}^{n}z_{t}^{(i)}(z_{t}^{(i)})^{\top})=\;$ $\sigma_{\min}^{+}(\sum\nolimits_{i=1}^{n}\hat{M}_{t}h_{t}^{(i)}(h_{t}^{(i)})^{\top}\hat{M}_{t}^{\top})$
=\;= $(\sigma_{\min}^{+}(\hat{M}_{t}))^{2}\sigma_{\min}\big{(}\sum\nolimits_{i=1}^{n}h_{t}^{(i)}(h_{t}^{(i)})^{\top}\big{)}$
=$\overset{(i)}{=}\;$ $\theta^{2}\cdot\Omega(n)=\Omega(\theta^{2}n),$
where $(i)$ holds with probability at least $1-p$, as long as $n\geq a\log(1/p)$ for some absolute constant $a>0$. This concentration is standard in analyzing linear regression (Wainwright, 2019).
=\;= $\mathcal{O}(\nu^{-1}mt^{5/2}(d_{y}+d_{u})^{2}d_{x}^{1/2}n^{-1/2}\log^{1/2}(1/p)),$
$\|\hat{z}_{t}-S_{t}z^{\ast}_{t}\|_{2}\leq\;$ $\|\hat{M}_{t}-S_{t}M^{\ast}_{t}\|_{2}\|h_{t}\|,$
from which we have that $\|\hat{z}_{t}-S_{t}z^{\ast}_{t}\|_{2}$ is sub-Gaussian with its mean and sub-Gaussian norm bounded by
$\mathcal{O}(\nu^{-1}mt^{3}(d_{y}+d_{u})^{5/2}d_{x}^{1/2}n^{-1/2}\log^{1/2}(1/p)).$
Consider
=$\overset{(i)}{=}\;$ $nd_{x}^{1/2}\log^{1/2}(n/p)\cdot\big{\|}\hat{z}_{t}^{(i)}-S_{t}(z^{\ast}_{t})^{(i)}\big{\|},$
where $(i)$ holds with probability $1-p$. Hence, there exists an absolute constant $c>0$, such that if $n\geq c\nu^{-6}m^{2}T^{6}(d_{y}+d_{u})^{5}d_{x}^{2}\log^{2}(n/p)$,
$\sigma_{\min}\big{(}\sum\nolimits_{i=1}^{n}\hat{z}_{t}^{(i)}(\hat{z}_{t}^{(i)})^{\top}\big{)}\geq\;$ $\sigma_{\min}\big{(}\sum\nolimits_{i=1}^{n}(z^{\ast}_{t})^{(i)}((z^{\ast}_{t})^{(i)})^{\top}\big{)}$
$\quad-\big{\|}\sum\nolimits_{i=1}^{n}\hat{z}_{t}^{(i)}(\hat{z}_{t}^{(i)})^{\top}-S_{t}(z^{\ast}_{t})^{(i)}((z^{\ast}_{t})^{(i)})^{\top}S_{t}^{\top}\big{\|}$
[…]
This is needed for the analysis of latent system identification in the next step.
Recovery of the latent dynamics. The latent system $(A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{T-1}$ is identified in Algorithm 3, using $(\hat{z}_{t}^{(i)})_{i=1,t=0}^{N,T}$ produced by Algorithm 2, by ordinary least squares. Recall from Proposition 1 that $z^{\ast}_{t+1}=A^{\ast}_{t}z^{\ast}_{t}+B^{\ast}_{t}u_{t}+L_{t+1}i_{t+1}$. With the transforms on $z^{\ast}_{t}$ and $z^{\ast}_{t+1}$, we have
and $(z^{\ast}_{t})^{\top}Q^{\ast}_{t}z^{\ast}_{t}=(S_{t}z^{\ast}_{t})^{\top}S_{t}Q^{\ast}_{t}S_{t}^{\top}S_{t}z^{\ast}_{t}$. Under control $u_{t}\sim\mathcal{N}(0,\sigma_{u}^{2}I_{d_{u}})$ for $10\leq t\leq T-$ 1, we know that $z^{\ast}_{t}$ is a zero-mean Gaussian random vector; so is $S_{t}z^{\ast}_{t}$. Let $\Sigma^{\ast}_{t}=\mathbb{E}[S_{t}z^{\ast}_{t}(z^{\ast}_{t})^{\top}S_{t}^{\top}]$ be its covariance.
For $10\leq t\leq\ell-$ 1, we need a bound for the estimation error of rank-deficient and noisy linear regression. By Lemma 5, there exists an absolute constant $c>0$, such that as long as $n\geq c(d_{x}+d_{u}+\log(1/p))$, with probability at least $1-p$,
=\;= $\mathcal{O}\big{(}(1+\beta^{-1})\|z_{t}-S_{t}z^{\ast}_{t}\|\log^{1/2}(n/p)+n^{-1/2}(d_{x}+d_{u}+\log(1/p))^{1/2}\big{)},$
which implies
$\|(\hat{A}_{t}-S_{t+1}A^{\ast}_{t}S_{t}^{\top})(\Sigma^{\ast}_{t})^{1/2}\|_{2}=\mathcal{O}((1+\beta^{-1})(t(d_{y}+d_{u}))^{3/2}d_{x}^{3/4}n^{-1/4}\log^{3/4}(n/p)),$
which is also a bound on $\|\hat{B}_{t}-S_{t+1}B^{\ast}_{t}\|_{2}$.
For $\ell\leq t\leq T-1$, by Lemma 6, with probability at least $1-p$,
=\;= $\mathcal{O}\big{(}(\nu^{-1}+\sigma_{u}^{-1})\big{(}\|\hat{z}_{t}-S_{t}z^{\ast}_{t}\|\log^{1/2}(n/p)+n^{-1/2}(d_{x}+d_{u}+\log(1/p))^{1/2}\big{)}\big{)},$
which implies
$\|\hat{A}_{t}-S_{t+1}A^{\ast}_{t}S_{t}^{\top}\|_{2}=\mathcal{O}\Big{(}(1+\nu^{-2})\big{(}mt^{3}(d_{y}+d_{u})^{5/2}d_{x}^{1/2}\log^{1/2}(1/p)\big{)}n^{-1/2}\Big{)},$
which is also a bound on $\|\hat{B}_{t}-S_{t+1}B^{\ast}_{t}\|_{2}$. We note that since the observation of $u_{t}$ is exact, a tighter bound on $\|\hat{B}_{t}-S_{t+1}B^{\ast}_{t}\|_{2}$ is possible; we keep the current bound since it does not affect the final bounds in Theorem 1.
By Assumption 3, $(Q^{\ast}_{t})_{t=0}^{\ell-1}$ and $Q^{\ast}_{T}$ are positive definite; they are identity matrices under the normalized parameterization. For $(Q^{\ast}_{t})_{t=\ell}^{T-1}$, which may not be positive definite, we identify them in SysId (Algorithm 3) by (3.5). By Lemma 2,
$\|\widetilde{Q}_{t}-S_{t}Q^{\ast}_{t}S_{t}^{\top}\|_{F}=\;$ $\mathcal{O}(md_{x}^{2}\log(n/p)\cdot mt^{3}(d_{y}+d_{u})^{5/2}d_{x}^{1/2}n^{-1/2}\log^{1/2}(1/p)$
$\quad+d_{x}^{2}md_{x}^{1/2}n^{-1/2}\log^{1/2}(1/p))$
=\;= $\mathcal{O}(m^{2}t^{3}(d_{y}+d_{u})^{5/2}d_{x}^{5/2}n^{-1/2}\log^{1/2}(1/p)),$
where we have used $\nu=\Omega(1)$. By construction, $\hat{Q}_{t}$ is the projection of $\widetilde{Q}_{t}$ onto the positive semidefinite cone with respect to the Frobenius norm (Boyd and Vandenberghe, 2004, Section 8.1.1). Since $S_{t}Q^{\ast}_{t}S_{t}^{\top}\succcurlyeq 0$,
RDE (2.4) proceeds backward. For $\ell\leq t\leq T-1$, $\|\hat{A}_{t}-S_{t+1}A^{\ast}_{t}S_{t}^{\top}\|_{2}$, $\|\hat{B}_{t}-S_{t+1}B^{\ast}_{t}\|_{2}$ and $\|\hat{Q}_{t}-S_{t}Q^{\ast}_{t}S_{t}^{\top}\|_{2}$ are all bounded by
$\mathcal{O}(m^{2}T^{3}(d_{y}+d_{u})^{5/2}d_{x}^{5/2}n^{-1/2}\log(1/p)^{1/2}).$
By Lemma 10 on certainty equivalent linear quadratic control, for $n\geq\mathrm{poly}(T,d_{x},d_{y},d_{u},\log(1/p))$, where hidden constants are dimension-free and depend polynomially on system parameters, the controller $(\hat{K}_{t})_{t=\ell}^{T-1}$ is $\epsilon$-optimal in system $((S_{t+1}A^{\ast}_{t}S_{t}^{-1},S_{t+1}B^{\ast}_{t},R^{\ast}_{t})_{t=\ell}^{T-1},(S_{t}Q^{\ast}_{t}S_{t}^{\top})_{t=\ell}^{T})$, for
ϵ=italic-$\epsilon=\;$ $\mathcal{O}((d_{x}\wedge d_{u})(d_{y}+d_{u})^{5}d_{x}^{5}m^{4}T^{7}\log(1/p)n^{-1}),$ (4.1)
that is, if
$n\geq cm^{4}(d_{y}+d_{u})^{5}d_{x}^{6}T^{7}\log(1/p)\epsilon^{-1},$
for a dimension-free constant $c>0$ that depends on system parameters.
For $10\leq t\leq\ell-$ 1, $(\hat{K}_{t})_{t=0}^{\ell-1}$ and $(K^{\ast}_{t})_{t=0}^{\ell-1}$ are optimal controllers in the $\ell$-step systems with terminal costs given by $\hat{P}_{\ell}$ and $P^{\ast}_{\ell}$, respectively, where $\hat{P}_{\ell}$ and $P^{\ast}_{\ell}$ are the solutions to RDE (2.4) in systems $((\hat{A}_{t},\hat{B}_{t},R^{\ast}_{t})_{t=0}^{T-1},(\hat{Q}_{t})_{t=0}^{T})$ and $((S_{t+1}A^{\ast}_{t}S_{t}^{-1},S_{t+1}B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(S_{t}Q^{\ast}_{t}S_{t}^{\top})_{t=0}^{T})$ at step $\ell$, respectively. By Lemma 10,
$\|\hat{P}_{\ell}-P^{\ast}_{\ell}\|_{2}=\mathcal{O}(m^{2}t^{3}(d_{y}+d_{u})^{5/2}d_{x}^{5/2}n^{-1/2}\log^{1/2}(1/p)),$
which is dominated by the $n^{-1/4}$ rate of $\|(\hat{A}_{t}-S_{t+1}A^{\ast}_{t}S_{t}^{\top})(\Sigma^{\ast}_{t})^{1/2}\|_{2}$ and $\|\hat{B}_{t}-S_{t+1}B^{\ast}_{t}\|_{2}$ for $10\leq t\leq\ell-$ 1. Then, by Lemma 9, $(\hat{K}_{t})_{t=0}^{\ell-1}$ is $\epsilon$-optimal in system $((S_{t+1}A^{\ast}_{t}S_{t}^{-1},S_{t+1}B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{\ell-1},(S_{t}Q^{\ast}_{t}S_{t}^{\top})_{t=0}^{\ell-1})$ with terminal cost matrix $P^{\ast}_{t}$, for
ϵ=italic-$\epsilon=\;$ $\mathcal{O}(a^{\ell}d_{x}(\ell(d_{y}+d_{u}))^{3/2}d_{x}^{3/4}n^{-1/4}\log^{3/4}(n/p)+d_{x}(d_{x}+d_{u}+\log(1/p))^{1/2}n^{-1/2}))$
=\;= $\mathcal{O}(d_{x}^{7/4}(\ell(d_{y}+d_{u}))^{3/2}a^{\ell}n^{-1/4}\log^{3/4}(n/p)),$ (4.2)
where dimension-free constant $a>0$ depends on the system parameters; that is, if
$n\geq a_{0}d_{x}^{7}(d_{y}+d_{u})^{6}\ell^{6}a_{1}^{\ell}\log^{3}(n/p)\epsilon^{-4},$
for some dimension-free constants $a_{0},a_{1}>0$ that depend on system parameters. The bound (4.2) is worse than (4.1), because for $10\leq t\leq\ell-$ 1, $z^{\ast}_{t}$ does not have full-rank covariance, and $S_{t+1}A^{\ast}_{t}S_{t}^{\top}$ is only recovered partially. Even with large enough data, linear regression has no guarantee for $\|\hat{A}_{t}-S_{t+1}A^{\ast}_{t}S_{t}^{\top}\|_{2}$ to be small; we do not know the controllability of $(\hat{A}_{t},\hat{B}_{t})_{t=0}^{\ell-1}$, not even its stabilizability. ∎
Next, we provide several key technical lemmas regarding the regression and factorization procedures used in our algorithm.

### v3

@@ 4 Theoretical guarantees
Theorem 1 below offers a finite-sample guarantee for our approach, confirming cost-driven state representation learning (Algorithm 1) as a viable path to solving LQG control.
@@
Given an unknown LQG control problem defined by (2.1) and (2.2), under Assumptions 1234 and 5, for a given $p\in(0,1)$, if we run Algorithm 1 for $n\geq\mathrm{poly}(T,d_{x},d_{y},d_{u},\log(1/p))$, then with probability at least 1p1-p, state representation function $(\hat{M}_{t})_{t=0}^{T}$ is $\mathrm{poly}(\ell,d_{x},d_{y},d_{u},\log(\ell/p))n^{-1/4}$-optimal in the first $\ell$ steps, and $\mathrm{poly}(\nu^{-1},T,d_{x},d_{y},d_{u},\log(T/p))n^{-1/2}$-optimal in the next $(T-\ell)$ steps. In addition, the overall output policy $\hat{\pi}=(\hat{M}_{t},\hat{K}_{t})_{t=0}^{T-1}$ satisties
$J(\hat{\pi})-J(\pi^{\ast})=\;$ $\mathrm{poly}(\beta^{-1},\ell,d_{x},d_{y},d_{u},\log(\ell/p))c^{\ell}n^{-1/4}$
$\qquad+\mathrm{poly}(\nu^{-1},m,T,d_{x},d_{y},d_{u},\log(nT/p))n^{-1},$
where $c>1$ is a dimension-free constant depending polynomially on $\ell$ and other problem parameters.
@@
Theorem 1 provides finite-sample guarantees for both the state representation function $(\hat{M}_{t})_{t=0}^{T}$ and the overall policy $\hat{\pi}=(\hat{M}_{t},\hat{K}_{t})_{t=0}^{T}$. From Theorem 1, we observe a separation of the sample complexities before and after time step $\ell$ for the state representation function, resulting from the loss of the full-rankness of $(\mathrm{Cov}(z^{\ast}_{t}))_{t=0}^{\ell-1}$ and $(M^{\ast}_{t})_{t=0}^{\ell-1}$. Specifically, quadratic regression guarantees that $\hat{N}_{t}$ converges to $N^{\ast}_{t}$ at a rate of $n^{-1/2}$ for all $0\leq t\leq T$ (Lemma 2). Before time step $\ell$, $\hat{M}_{t}$ suffers a square root decay of the rate $n^{-1/4}$ because $M^{\ast}_{t}$ may not have rank $d_{x}$ (Lemma 4). Since $(\hat{z}_{t})_{t=0}^{\ell-1}$ may not have full-rank covariances, $(A^{\ast}_{t})_{t=0}^{\ell-1}$ are only recovered partially. As a result, $(\hat{K}_{t})_{t=0}^{\ell-1}$ may not stabilize $(A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{\ell-1}$. This issue poses a significant challenge for the analysis and significantly worsens dependence on $\ell$ in the policy suboptimality gap (Lemma 7), which implies that if $n$ is not large enough, the policy may be inferior to zero control in the first $\ell$ steps, as system $(A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{\ell-1}$ is uniformly exponential stable (Assumption 1), while zero control has a suboptimality gap linear in $\ell$. After time step $\ell$, $\hat{M}_{t}$ retains the $n^{-1/2}$ sample complexity, from which the same order of sample complexity guarantee for $(\hat{A}_{t},\hat{B}_{t})$ follows, resulting in the $\mathrm{poly}(T)n^{-1}$ term in the policy suboptimality gap.
Next, we provide a key proposition and several technical lemmas for analyzing our algorithm, before we present the proof of Theorem 1, which contains the exact polynomial orders of the finite-sample guarantees.
4.1 Proposition on multi-step cumulative costs
The following proposition establishes the relationship between the multi-step cumulative costs and the state estimates by the Kalman filter.
Proposition 3.
Let $(z^{\ast\prime}_{t})_{t=0}^{T}$ be the state estimates by the Kalman filter under the normalized parameterization. If we apply $u_{t}\sim\mathcal{N}(0,\sigma_{u}^{2}I)$ for all $0\leq t\leq T-1$, then for $0\leq t\leq T$,
where $k=1$ for $0\leq t\leq\ell-1$ and $k=m\wedge(T-t+1)$ for $\ell\leq t\leq T$, $\overline{b}_{t}=\mathcal{O}(k)$, and $\overline{e}_{t}$ is a zero-mean subexponential random variable with $\|\overline{e}_{t}\|_{\psi_{1}}=\mathcal{O}(kd_{x}^{1/2})$.
@@ 4.1 Proposition on multi-step cumulative costs
$\overline{c}_{t}=\;$ $\sum\nolimits_{\tau=t}^{t+k-1}c_{\tau}$
$=\;$ $\sum\nolimits_{\tau=t}^{t+k-1}\big{(}\|\Phi^{\prime}_{\tau,t}z^{\ast\prime}_{t}+j^{\prime}_{\tau,t}\|_{Q^{\ast\prime}_{\tau}}^{2}+\|u_{\tau}\|_{R^{\ast}_{\tau}}^{2}+b^{\prime}_{\tau}+e^{\prime}_{\tau}\big{)}$
$=\;$ $(z^{\ast\prime}_{t})^{\top}\Big{(}\sum\nolimits_{\tau=t}^{t+k-1}(\Phi^{\prime}_{\tau,t})^{\top}Q^{\ast\prime}_{t}\Phi^{\prime}_{\tau,t}\Big{)}z^{\ast\prime}_{t}+\sum\nolimits_{\tau=t}^{t+k-1}\|u_{\tau}\|_{R^{\ast}_{\tau}}^{2}$
$=\;$ $\|z^{\ast\prime}_{t}\|^{2}+\sum\nolimits_{\tau=t}^{t+k-1}\|u_{\tau}\|_{R^{\ast}_{\tau}}^{2}+\overline{b}_{t}+\overline{e}_{t},$